Calculate the pH of uneous solution of 0.234 M sodium sulfite. Retry Entire Group 9 more group attempts remaining
Answer #1

Potassium sulfite in K2SO3 It is a wrong erecevolyte and it dimeciates completely in water to art and so 2 - Soz is a conjugaSo} + H2O 2 HSO3 + oH I 0.234 c -X +x. ta th E 0.234-2 Kbi = [1sozlequillibrium Cow Jewillibrium 0.234-2 0.159x105 - * 0.234-H₂O ² HSO3 + 0.193x103 H2SO3 + O OH- 0.193x163 I c ta + x -X [ 0.193x10²x - - 0.193x183+a Neglect (0.193x16? Kb = considering1. Total [om-] = (0.193x163 + 0.14 x1õl2 ) m = 0.193x173 m pole = -log(OH-] = -log(0-193x163 ) = 3-714 pH = 14 - pOH 14-3.714